# Quantum Byzantine Multiple Access Channels

### v1

@@ Abstract
In communication theory, attacks like eavesdropping or jamming are typically assumed to occur at the channel level, while communication parties are expected to follow established protocols. But what happens if one of the parties turns malicious? In this work, we investigate a compelling scenario: a multiple-access channel with two transmitters and one receiver, where one transmitter deviates from the protocol and acts dishonestly. To address this challenge, we introduce the Byzantine multiple-access classical-quantum channel and derive an achievable communication rate for this adversarial setting.
@@ I Introduction
As a result, many of the sophisticated techniques from [29] cannot be directly extended to quantum Byzantine multiple-access channels. Similarly, the identification technique proposed in [30] is not applicable: while a sequence of classical symbols can be easily verified against a set of expected legal channel outputs, a “false” quantum state may still have a nonzero probability of being accepted by the decoding operators for legitimate messages. To overcome these challenges, we must develop a novel coding approach.
@@ II Preliminaries
Notice that when we choose an “optimal” set $\Phi:=\{\rho_{x}:x\in X\}$ (in the sense that it maximizes the quantum mutual information and the Holevo quantity), we have (cf. [21])
@@ III Message Transmission over Quantum Byzantine Multiple Access Channel
III-A Necessity of Decoding Order
@@
However, in the Byzantine multiple-access classical-quantum channel scenario, we cannot apply the Tender Operator trick, since we cannot guarantee that the first decoding operator will decode the first message with a sufficiently high probability, as the first sender may indeed be the adversarial one. Therefore, in this scenario, it is possible that if the first sender is adversarial, the first decoding operator may alter the channel outputs in such a way that the second message becomes undecodable. The following simple example illustrates this point:
@@
When the receiver, unaware that sender $2$ is adversarial, uses $D_{0}^{(2)}$ and $D_{1}^{(2)}$ to measure sender $2$’s message, there is a $50\%$ chance that the receiver obtains $m_{0}^{2}$, with the resulting channel output after the measurement given by:
@@
The same scenario occurs when sender $1$ transmits $m_{1}^{1}$ (i.e., when $1$ is sent through the channel) and sender $2$ jams the channel by sending $2$. In this case, the “worse order”, where the adversarial message is decoded first, results in the second code becoming ineffective due to the interference of the first POVM decoding operator. Therefore, in case $2c$, decoding the message of the adversarial sender first renders the code unreliable.
@@ III-B Main Result
$R_{1}\leq\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),$
$R_{2}\leq\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),$
where $E_{1}B_{x^{n}_{1}}$ represents the outcome of $C_{E_{1}}\circ\mathtt{W}((x^{n}_{1}(i),\cdot))$.
@@
Thus, after the measurement of sender 1’s message, when the receiver begins to decode sender 2’s message, the message transmission rate cannot exceed the arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ within the subspace defined by $D_{m_{1}}^{\gamma_{1},\gamma_{2}}$. The arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ into this subspace is equal to
$\min_{x^{n}_{1}}\chi(p_{2},\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2})))^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}).$
@@
$\sum_{m_{1}}\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)\min_{x^{n}_{1}}\chi\left(p_{2},\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\right),$
@@
The capacity region of $\mathtt{W}$ is included in the following:
$R_{1}$ $\leq\min\left(\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),\max_{p_{1}}\min_{x^{n}_{3}}I(p_{1},B_{x^{n}_{3}})\right)$ (1)
$R_{2}$ $\leq\min\left(\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),\max_{p_{2}}\min_{x^{n}_{3}}I(p_{2},B_{x^{n}_{3}})\right)$ (2)
$R_{3}$ $\leq\min\left\{\max_{p_{3}}\min_{x^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\}$ (3)
where for $i=1,2,3$, $B_{x^{n}_{i}}$ represents the channel outcome when sender $i$ sends $x^{n}_{i}$ as a jamming attack.
@@
The achievable rates for the messages of sender 1 and sender 2 follow from Theorem III.1. Now, consider the situation where the receiver has already performed measurements using the codes $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$, which are permutations of the code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, and $\{x^{\gamma_{2}}_{2}(m_{2}),D_{m_{2}}^{\gamma_{2}}:m_{2}\in M_{2}\}$, which are permutations of the code $\{x^{n}_{2}(m_{2}),D_{m_{2}}:m_{2}\in M_{2}\}$. The receiver now aims to decode the message of sender 3.
@@
$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))=I(p_{3},E_{1}E_{2}B_{x^{n}_{2}}|p_{1}).$
Notice that when the code for the legitimate sender is reliable, then, by the Tender Operator Lemma, only the jammer’s decoding operator affects the channel output. Therefore, by [5] and [14], and if we set
$|M_{3}|$ $<\max_{p_{3}}\min_{t\in\Theta}\chi(p_{3},\mathbf{W}(p_{3},t))$
$=\min\left(\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{1})),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))\right)$
$=\min\left\{\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\},$
@@
The random capacity region of $\mathtt{W}$ is included in
$R_{i}\leq\min\Bigl{(}\max_{p_{i}}\min_{x^{n}_{j}:j<i}I(p_{i},E_{j}B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\}),$ (5)
$\max_{p_{i}}\min_{x^{n}_{j}:j>i}I(p_{i},B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\})\Bigr{)}\text{ .}$ (6)
@@ IV Conclusion and Future Directions
A general upper bound for the rate region with more than three senders remains an open problem. Furthermore, an important challenge lies in applying advanced classical techniques, such as adversarial sender identification from [29] and [30], to quantum channels. The main difficulty, as mentioned earlier, is that for quantum arbitrarily varying channels, only random coding techniques are currently known.
Acknowledgments
The authors acknowledge the support of the German Federal Ministry of Education and Research (BMBF) under several national initiatives: the 6G Communication Systems initiative through the research hub 6G-life (grant 16KISK263), the QuaPhySI (Quantum Physical Layer Service Integration) initiative (grant 16KIS2234), the “QTOK - Quantum Tokens for Secure Authentication in Theory and Practice” initiative (grant 16KISQ038), the QR.N initiative (grant 16KIS2196), the QTREX initiative (grant 16KISR038), the QUIET initiative (grant 16KISQ0170), and the QD-CAMNetz initiative (grant 16ISQ169).

### v2

@@ Abstract
In communication theory, attacks like eavesdropping or jamming are typically assumed to occur at the channel level, while communication parties are expected to follow established protocols. But what happens if one of the parties turns malicious? In this work, we investigate a compelling scenario: a multiple-access channel with two transmitters and one receiver, where one transmitter deviates from the protocol and acts dishonestly. To address this challenge, we introduce the Byzantine multiple-access classical-quantum channel and derive the communication region of rate for this adversarial setting.
@@ I Introduction
As a result, many of the sophisticated techniques from [29] cannot be directly extended to quantum Byzantine multiple-access channels. Similarly, the identification technique proposed in [30] is not applicable: while a sequence of classical symbols can be easily verified against a set of expected legal channel outputs, a "false" quantum state may still have a nonzero probability of being accepted by the decoding operators for legitimate messages. To overcome these challenges, we must develop a novel coding approach.
@@ II Preliminaries
Notice that when we choose an "optimal" set $\Phi:=\{\rho_{x}:x\in X\}$ (in the sense that it maximizes the quantum mutual information and the Holevo quantity), we have (cf. [21])
@@
For a probability distribution $p$ on $\Theta$, we denote $\mathbf{W}(\cdot,p):=\sum_{t}\mathbf{W}(\cdot,t)$.
@@ III Message Transmission over Quantum Byzantine Multiple Access Channel
III-A Impact of Decoding Order
@@
However, in the Byzantine multiple-access classical-quantum channel scenario, we cannot apply the Tender Operator trick, since we cannot guarantee that the first decoding operator will decode the first message with a sufficiently high probability, as the first sender may indeed be the adversarial one. Therefore, in this scenario, it is possible that if the first sender is adversarial, the first decoding operator may alter the channel outputs in such a way that the second message becomes undecodable. The following simple one-shot example illustrates this point:
@@
When the receiver, unaware that sender $2$ is adversarial, uses $D_{0}^{(2)}$ and $D_{1}^{(2)}$ to measure sender $2$’s message, there is a $50\%$ chance that the receiver obtains either $m_{0}^{2}$ or $m_{1}^{2}$. When he obtains $m_{0}^{2}$, then the resulting channel output after the measurement is given by:
@@
The same scenario occurs when sender $1$ transmits $m_{1}^{1}$ (i.e., when $1$ is sent through the channel) and sender $2$ jams the channel by sending $2$. In this case, the "worse order", where the adversarial message is decoded first, results in the second code becoming ineffective due to the interference of the first POVM decoding operator. Therefore, in case $2c$, decoding the message of the adversarial sender first renders the code unreliable.
@@ III-B Main Result
$R_{1}\leq\max_{p_{1}}\min_{p_{2}}I(p_{1},B),$
$R_{2}\leq\max_{p_{2}}\min_{p_{1}}I(p_{2},E_{1}B),$
where $B$ represents the outcome of $\mathtt{W}((p_{1},p_{2})$. and $E_{1}B$ represents the outcome of $C_{E_{1}}\circ\mathtt{W}(p_{1},p_{2})$.
@@
Thus, after the measurement of sender 1’s message, when the receiver begins to decode sender 2’s message, the message transmission rate cannot exceed the arbitrarily varying channel capacity of $\mathtt{W}(p_{1},p_{2})$ within the subspace defined by $D_{m_{1}}^{\gamma_{1},\gamma_{2}}$. The arbitrarily varying channel capacity of $\mathtt{W}((p_{1},\cdot))$ into this subspace is equal to
$\min_{x^{n}_{1}}\chi(p_{2},\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((p_{1},x^{n}_{2})))^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{p_{1},p_{2}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}).$
@@
$\sum_{m_{1}}\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)\min_{x^{n}_{1}}\chi\left(p_{2},\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\right),$
@@
The capacity region of $\mathtt{W}$ is the following:
$R_{1}$ $\leq\min\left(\max_{p_{1}}\min_{p_{2}}I(p_{1},B),\max_{p_{1}}\min_{p_{3}}I(p_{1},B)\right)$ (1)
$R_{2}$ $\leq\min\left(\max_{p_{2}}\min_{p_{1}}I(p_{2},E_{1}B),\max_{p_{2}}\min_{p_{3}}I(p_{2},B)\right)$ (2)
$R_{3}$ $\leq\min\left\{\max_{p_{3}}\min_{p_{1}}I(p_{3},E_{1}B|p_{2}),\max_{p_{3}}\min_{p_{2}}I(p_{3},E_{2}B|p_{1})\right\}$ (3)
@@
The capacity region for the messages of sender 1 and sender 2 follow from Theorem III.1. Now, consider the situation where the receiver has already performed measurements using the codes $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$, which are permutations of the code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, and $\{x^{\gamma_{2}}_{2}(m_{2}),D_{m_{2}}^{\gamma_{2}}:m_{2}\in M_{2}\}$, which are permutations of the code $\{x^{n}_{2}(m_{2}),D_{m_{2}}:m_{2}\in M_{2}\}$. The receiver now aims to decode the message of sender 3.
@@
$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))=I(p_{3},E_{1}E_{2}B|p_{1}).$
Denote the set of probability distributions on $\Theta$ by $P(\Theta)$. Notice that when the code for the legitimate sender is reliable, then, by the Tender Operator Lemma, only the jammer’s decoding operator affects the channel output. Therefore, by [5] and [14], and if we set
$|M_{3}|$ $<\max_{p_{3}}\min_{p_{t}\in P(\Theta)}\chi(p_{3},\mathbf{W}(p_{3},p_{t}))$
$=\min\left(\max_{p_{3}}\min_{p_{1}}\chi(p_{3},\mathbf{W}(p_{3},p_{1})),\max_{p_{3}}\min_{p_{2}}\chi(p_{3},\mathbf{W}(p_{3},p_{2}))\right)$
$=\min\left\{\max_{p_{3}}\min_{p_{1}}I(p_{3},E_{1}B|p_{2}),\max_{p_{3}}\min_{p_{2}}I(p_{3},E_{2}B|p_{1})\right\},$
@@
For the converse for the message of sender 3, we may, without loss of generality, assume that the receiver uses a separable decoding strategy, as stated in the proof of Thoerem III.1 . We consider now the case when sender 3 sends $x^{\gamma_{3}}_{3}(m_{3})$ through the channel. When sender 1 is adversarial and sends $x^{n}_{1}$ as jamming attack, bthen by the Tender Operator Lemma, the channel output will be sufficiently closed to
$\sum_{m_{2}}p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))\,\sqrt{D_{m_{1}}}$
with probability $\mathrm{tr}\biggl{(}\sum_{m_{2}}p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))$ $\sqrt{D_{m_{1}}}$ $\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))$ $\sqrt{D_{m_{1}}}\biggr{)}$. Thus, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{1}))$, When sender 2 is adversarial and sends $x^{n}_{2}$ as jamming attack, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{2}))$. By [5] and [14], the rate for the message of sender 3 cannot exceed
$\max_{p_{3}}\min_{p_{1}}\chi(p_{3},\mathbf{W}(p_{3},p_{1}))$
or
$\max_{p_{3}}\min_{p_{2}}\chi(p_{3},\mathbf{W}(p_{3},p_{2}))\text{ .}$
This gives us the upperbound.
@@
The random capacity region of $\mathtt{W}$ is
$R_{i}\leq\min\Bigl{(}\max_{p_{i}}\min_{p_{j}:j<i}I(p_{i},E_{j}B|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\}),$ (5)
$\max_{p_{i}}\min_{p_{j}:j>i}I(p_{i},B|\{p_{h}:h\in\{1,\dots,i-1\}\})\Bigr{)}\text{ .}$ (6)
@@
[…]
$D_{m_{i}}=\bigotimes_{m_{1},\cdots,m_{i-1}}D_{m_{i}}^{(m_{1},\cdots,m_{i-1})}$
for some operators $\{D_{m_{i}}^{(m_{1},\cdots,m_{i-1})}:m_{1},\cdots,m_{i-1}\}$.
@@
The converse can be shown in a similar as the proof of the converse of Theorem III.2. We consider now the case when sender i sends $x^{\gamma_{i}}_{i}(m_{i})$ through the channel. Let us define an arbitrarily varying classical-quantum channel
$\{\mathsf{W}(\cdot,t):t\in\Theta_{i}\}:X_{i}\rightarrow\mathcal{S}(A)$
denoted by
$\mathsf{W}(x^{n}_{i},x^{n}_{j})$
$\quad\cdot\sqrt{D_{m_{j}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{k}(m_{k})))\sqrt{D_{m_{j}}}$
for $j<i$, and
$\mathsf{W}(x^{n}_{i},x^{n}_{j})$ $=\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i})$
$\quad\cdot\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
for $j>i$. When sender j is adversarial and sends $x^{n}_{j}$ as jamming attack, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathsf{W}(x^{\gamma_{i}}_{i}(m_{i}),x^{n}_{j})$. By [5] and [14], the rate for the message of sender i cannot exceed
$\min_{p_{j}}I(p_{i},E_{j}B|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\})$
when $j$ $<i$, and
$\min_{p_{j}}I(p_{i},B|\{p_{h}:h\in\{1,\dots,i-1\}\})$
when $j$ $<i$. This gives us the upperbound.
@@ IV Conclusion and Future Directions
An important challenge lies in applying advanced classical techniques, such as adversarial sender identification from [29] and [30], to quantum channels. The main difficulty, as mentioned earlier, is that for quantum arbitrarily varying channels, only random coding techniques are currently known.
Acknowledgments.
The authors acknowledge the support of the German Federal Ministry of Education and Research (BMBF) under several national initiatives: the 6G Communication Systems initiative through the research hub 6G-life (grant 16KISK263), the QuaPhySI (Quantum Physical Layer Service Integration) initiative (grant 16KIS2234), the "QTOK - Quantum Tokens for Secure Authentication in Theory and Practice" initiative (grant 16KISQ038), the QR.N initiative (grant 16KIS2196), the QTREX initiative (grant 16KISR038), the QUIET initiative (grant 16KISQ0170), and the QD-CAMNetz initiative (grant 16ISQ169).